OpenStudy (anonymous):

For which equation does the sum of the roots equal -3 and the product of the roots equal 2? Answer a. 2x2 - 6x + 4 = 0 b. x2 + 2x - 3 = 0 c. x2 - 3x + 2 = 0 d. 2x2 + 6x + 4 = 0

OpenStudy (cwrw238):

for equation ax^2 + bx + c = 0 sum of roots = -b/a and product of roots = c/a

OpenStudy (anonymous):

i got c what did u get

OpenStudy (cwrw238):

you r correct